OpenStudy (anonymous):

4[-5x+7-3(x+1)]=-5x+4 i need step by step help

OpenStudy (radar):

On the first line that should be a -3 4[-5x+7-3x-3] -20x+28-12x-12 -32x+16=-5x+4 -27x=-12 x=-12/-27 = 4/9

OpenStudy (radar):

Error removed.

OpenStudy (radar):

Any questions???

OpenStudy (radar):

Good luck with it